# MS-XXXX-XXXX.XX

### v1

@@
We use valid inequalities (cuts) of the binary integer program for winner determination in a combinatorial auction (CA) as “artificial items” that can be interpreted intuitively and priced to generate Artificial Walrasian Equilibria. While the lack of an integer programming gap is sufficient to guarantee a Walrasian equilibrium, we show that it does not guarantee a “price-match equilibrium” (PME), a refinement that we introduce, in which prices are justified by an iso-revenue outcome for any hypothetical removal of a single bidder. We prove the existence of PME for any CA and characterize their economic properties and computation. We implement minimally artificial PME rules and compare them with other prominent CA payment rules in the literature.
@@ 2 Background and Related Literature
Approximation to competitive equilibrium is also used in allocation markets with fictional money by Budish (2011), a format that has proved useful in practice for seats in academic courses and other situations of heterogeneous preferences and high-demand. More recent literature along this thread is given by Nguyen and Vohra (2022) who investigate social-approximate equilibria when no WE exists, providing both approximation techniques via relaxation and preference restrictions that result in provable bounds on the relaxation error. Taken as a whole, this literature stream explores a number of ways to approach practical market design by approximating WE along various dimensions when they are not guaranteed to exist, as in the current work.
@@
When the coalitional valuation function (WDP considered over any subset of bidders) is submodular the minimal NWE and VCG outcomes are guaranteed to be equivalent (de Vries et al. 2007), and therefore the revenue comparison (1) above flattens, yielding strict equality among the NWE, MAP, MRC, and VCG solution concepts. (Coalitional submodularity holds when any bidder always adds weakly more value to a set of bidders than any of its supersets.) Under such a strict assumption on preferences, the auction of that paper can be run to deliver the same outcome we prescribe here, VCG with a decomposition into linear item prices (a WE). Here, we make no such assumption but additionally provide distinct MAP prices in cases where coalitional submodularity fails.
@@
Also related to the current work are others that implicitly introduce cuts and price them. Among these, Milgrom and Watt (2022) find market outcomes that approximate natural WE when none exist. They do not make the distinction of desiring PMEs as introduced here, nor do they explore the concept of cuts as artificial items directly. Like O’Neill et al. (2005), they have a single markup offset which can be interpreted in the language of the current paper as a single cut to the WDP-LP based on prices, as if a tax on auction spending. The combinatorial cuts of the current paper attempt instead to generate artificial structures that are as simple or explainable as possible, in that we look for cuts reflecting simple relationships among bids that can be represented as integer knapsack constraints, with smaller constant terms favored as simpler and more explainable.
@@
Even more closely related is the recent work of Lahaie and Lubin (2019) who introduce the adaptive (endogenous) generation of price terms in a dynamic (tâtonnement) CA. They focus on CAs with unique items (supply of one item for each commodity) and “monomial” price terms that each correspond to a subset of items, allowing for the pricing of certain synergies directly, with the goal of terminating at WE prices. From the perspective of the current paper, each monomial price term can be viewed as a clique cut (i.e., a fictional item for which there is exactly one copy), though it can also be shown that not all clique cuts are implicitly considered. (For example, Example  4.1 below uses a clique cut that does not correspond to a unique subset or monomial term in that paper.) They are able to show computationally that these cuts are highly effective in generating WEs by augmenting the space of prices under consideration, complementing the perspectives offered here. Here we also show a framework that adds pricing terms, but we move to the more general multi-unit CAs and explicit search over a broader set of possible cuts.
@@
All AWE (and hence PME) prices naturally produce core-selecting auctions, in which no coalition of bidders could propose a mutually (weakly) preferred outcome with higher total revenue. Core-selecting auctions are also described in Day and Milgrom (2008), with the precise selection of core points discussed in Day and Cramton (2012) and Bünz et al. (2022). We use the ideas of quadratic minimization as a tie-breaking rule as in the latter papers, but emphasize that here, quadratic tie-breaking is over item prices instead of personal total payments, an arguably better approach. Here we also show that all core-selecting points can be reached as an AWE via just one valid cut, which allows us to easily select the core point we want: PME prices with a minimal amount of artificial price adjustment, decomposed into a collection of more interpretable cuts.
@@ 3 Model Preliminaries
In a combinatorial auction (CA), a single seller offers $J$ different types of indivisible items for sale to a set of $I$ bidders, denoted =$\mathcal{I}=\{1,2,...i,...I\}$. The set of item types is denoted $\mathcal{J}=\{1,2,...j,...J\}$, and each item $j$ has a supply of $c_{j}\geq 1$ identical copies available to be purchased. CAs with unique items, i.e., those with $c_{j}=1$ for all $j$, are often studied in the literature; here we allow for the more general case. Column vector $c$ will denote the complete supply vector with elements $c_{j}$.
@@ 3.1 Winner Determination
The linear relaxation of WDP replaces (4) with $\bm{0}\leq x\leq\bm{1}$, and will be called WDP-LP. Because a solution to WDP-LP is potentially fractional, we write $w^{f}(\mathcal{I},\mathcal{J},\mathcal{K})$ for its objective function. Similarly, $x^{f}$ is a solution to WDP-LP, used when it is different from $x^{*}$.
@@ 3.3 Computing Dual Prices
In this paper, we explore finding AWEs, which will act like standard WEs when artificially added items are treated as real items. Accordingly, if we assume $\bm{A}$ contains a final set of added cuts, WDP-LP will solve to integer optimality, and dual prices form a WE per the main result of Bikhchandani and Mamer (1997). The dual of WDP-LP is:
@@ 4.1 Motivation and examples
which combine to yield $p_{A}+p_{B}+p_{C}\geq 15$, a violation of IR with $15>12$. It is easy to show that the associated WDP has a valid cut, an inequality that maintains all integer solutions but may exclude fractional solutions to WDP-LP:
@@ 4.2 Price-Match Equilibria for CAs
Adomavicius and Gupta (2005) define a winning level of a bundle $a$ at any instant in a continuous efficient CA as the amount needed for a new bid on $a$ to just become winning, given the bids submitted up until that point in time. They show that it can be computed in our notation (before the new bid has been added to $\mathcal{K}$) as $w(\mathcal{I},\mathcal{J},\mathcal{K})-w(\mathcal{I},\mathcal{J}\setminus a,\mathcal{K})$, where $\mathcal{J}\setminus a$ represents the same item classes but with supply vector reduced to $c-a$.
@@
Indeed, in Example  4.1 , without the cut, the first three bids would be “half accepted” by solution $x^{f}=[0.5,0.5,0.5,0]^{T}$ to WDP-LP. In doing so, it’s as if the three bids are allowed to represent $15$ units of value, when really they only generate $10$ units of value integrally. In Example 4.2 (if a cut is not applied), the raised bid “hides” this non-convex over-influence of these three bids, allowing them to drive prices up beyond their ability to match those prices in any integral solution.
To further refine the winning-level notion for the current context (i.e., a multi-unit, sealed-bid CA with linear prices and multi-minded bidders), consider the following definition:
@@
Directly from the definition of PME, the second-price auction for a single item has a PME outcome. Similarly, the Generalized Second-Price Auction, in which the winner of the $n$th slot pays the amount of the $n+1$st bid (famous for its use in Google’s ad-word auctions), prescribes a PME outcome: if any bidder were to leave, each bidder in a worse-ranked slot could be moved up one slot without violating IR. Considering the more general CA environment, per de Vries et al. (2007), when coalitional submodularity holds, VCG is implementable as a NWE. Together with Corollary 6.9 below, we have:
@@
Coalitional submodularity is sufficient to guarantee min-NWE = PME = MRC = VCG.
@@ 5 Comparison to Min-Revenue Core-Selecting and VCG Mechanisms
The proof sketch is that core-selecting constraints as in Equation 9 can be derived by simply adding the relevant envy-freeness constraints (5) indexed by winning bids in the $w(\mathcal{C},\mathcal{J},\mathcal{K^{\mathcal{C}}})$ outcome. The converse does not directly hold: not all core outcomes can be reached via the currently available items prices. (This will be addressed later by Theorem 6.3, however.) To demonstrate this important difference between MAP prices and the VCG and MRC benchmarks from the literature, consider the following numerical example, adapted from Beck (2013):
@@
Figure 2: The core of Example 5.2 in the $\bm{\rho_{1}}$,$\bm{\rho_{2}}$ plane. The lighter region consists of NWE payment outcomes while the AWE $\bm{=}$ PME outcomes form the gray shaded triangle. The thicker line segment intersection of these two regions is the image of MAP prices on the core. WDP-quad-dual selects MAP (red point) from among the PME points. The pay-as-bid and MRC $\bm{=}$ VCG points are also labeled.
@@
$\rho^{MRC}_{1}=5\text{ for A}$ $\rho^{MRC}_{2}=5\text{ for B}$ when selling $\text{when selling }\{A,B\}$
$\rho^{MRC}_{1}=10\text{ for C}$ when selling $\text{when selling }\{C\}$
$\rho^{MRC}_{1}=10\text{ for AC}$ $\rho^{MRC}_{2}=0\text{ for B}$ when selling $\text{when selling }\{A,B,C\}$
and observe that total revenue goes down drastically (from $20$ to $10$) when combining the $\{A,B\}$ auction with that of an “unrelated good” $\{C\}$ into a single auction. Beck (2013) notes that this shows a similar “low-revenue” problem as VCG, observing that total payments collected by VCG are $10$ whether these two auctions are conducted separately or together. With this example as motivation, we observe the unrelated goods problem vanish for MAP prices as proposed here.
@@ 6 Generating Cuts for WE and PME
We first consider generating a single valid cut. Let $\alpha_{0}$ be the right-hand side of a new cut, with cut coefficients $\alpha_{k}$ for each $k\in\mathcal{K}$, the latter collectively forming the vector $\alpha$, together forming a cut:
@@
which must hold for any WDP-feasible allocation $x\in\{0,1\}^{K}$. These $\alpha$ terms will be variables during the cut generation process and then later treated as constants in the final price determination, functioning exactly as all other items in $\mathcal{J}$, and concatenated as a new row of the WDP matrix $\bm{A}$. This new cut will thus be treated as an artificial item in the auction. Eventually, we prefer a collection of artificial items with each $(\alpha,\alpha_{0})\in\mathbb{Z}^{K+1}_{\geq 0}$ for interpretability, but for intermediate steps in our reasoning and algorithms we will also consider “unscaled” or continuous valid cuts, which we denote with a bar-notation: $(\bar{\alpha},\bar{\alpha}_{0})\in\mathbb{R}^{K+1}_{\geq 0}$.
Any newly added artificial item must be fully demanded at our fixed WDP solution $x^{*}$, i.e., $\alpha x^{*}=\alpha_{0}$. That is every copy of an artificial item must be purchased by the winning bids. If not, WE fails by definition, and pricing an excess-supply item may violate the core property, even when envy-freeness constraints (5) are satisfied. Because the preponderance of previous work on cut generation was used to solve IPs like WDP, this assumption that we know $x^{*}$ and constrain that $\alpha x^{*}=\alpha_{0}$ when looking for cuts is peculiar to the current line of research, to our knowledge.

### v2

@@
We use valid inequalities (cuts) of the binary integer program for winner determination in a combinatorial auction (CA) as “artificial items” that can be interpreted intuitively and priced to generate Artificial Walrasian Equilibria. We thus provide a method for converting a CA problem that admits only non-anonymous, nonlinear bundle prices into one that admits anonymous linear prices over the augmented item space, forestalling ex-post bidder complaints about opaque and strongly discriminatory pricing. To this end, we introduce a refinement of the Walrasian equilibrium which we call a “price-match equilibrium” (PME) in which all prices are justified by providing an iso-revenue reallocation for the hypothetical removal of any single bidder. We prove the existence of PME for any CA and characterize their economic properties and computation. We implement minimally artificial PME rules and compare them with other prominent CA payment rules in the literature.
@@ 2 Background and Related Literature
Approximation to competitive equilibrium is also used in allocation markets with fictional money by Budish (2011), a format that has proved useful in practice for seats in academic courses and other situations of heterogeneous preferences and high-demand. More recent work along this line is given by Nguyen and Vohra (2022) who investigate social-approximate equilibria when no WE exists, providing both approximation techniques via relaxation and preference restrictions that result in provable bounds on the relaxation error. Taken as a whole, this literature stream explores a number of ways to approach practical market design by approximating WE along various dimensions when they are not guaranteed to exist, as in the current work.
@@
When each bidder is interested in consuming at most one item (a.k.a., unit demand) the minimal NWE and VCG outcomes are guaranteed to be equivalent (Leonard 1983), and therefore the revenue comparison (1) above flattens, yielding strict equality among the minimal NWE, MAP, MRC, and VCG solution concepts. Under such a strict assumption on preferences, several known auction mechanisms (e.g., primal-dual auctions as in de Vries et al. (2007), etc.) could be run to deliver the same outcome we prescribe here: VCG with a decomposition into linear-anonymous WE item prices. Here, we make no such preference assumptions and instead provide distinct MAP prices for cases in which unit demand and even less strict conditions like gross substitutes and coalitional submodularity (de Vries et al. 2007) fail.
Bikhchandani and Ostroy (2002) show that VCG is a WE if and only if a “buyers are substitutes” condition holds. They explored combinatorial exchanges (i.e., with potentially several sellers in addition to several buyers), and consider equilibria supported by potentially non-linear and non-anonymous prices. Under general preferences for CAs (i.e., with one seller under standard assumptions), they show the existence of instances for which only non-linear and non-anonymous pricing equilibria exist. Here, we show that if one is willing to accept an extended notion of linear prices on carefully selected artificial items, then linear anonymous prices are possible whenever the core is non-empty, including all such CAs.
The introduction of artificial items results in two distinct notions of anonymous prices: (1) as in Bikhchandani and Ostroy (2002), prices are anonymous if any two bidders see the same total price for the same bundle of natural items; (2) herein, prices are anonymous if any two bidders see the same total price for the same bundle of natural and artificial items. Note that (2) does not imply (1); in Example 6.9 (below) bidder 1 sees a price for bundle $\{B,C\}$ of 12, while bidder 2 sees a price of 8, violating (1). Yet, (2) holds because only bidder 1 would be required to buy artificial items priced at 4 along with $\{B,C\}$, based on the combinatorial structure, as will be explained below. Accordingly, interpreting artificial items as licenses the seller could have offered, all bidders face the same traditionally anonymous price vector (i.e. according to definition (1) above) in this augmented setting, which is precisely what we mean by definition (2). Mirroring the Bikhchandani and Ostroy (2002) characterization of their non-linear non-anonymous equilibria as “filling the core” when there is only one seller, we show that the linear anonymous prices (under definition (2) above) that we introduce also fill the core in Theorem 6.2.
@@
Also related to the current work are others that implicitly introduce cuts and price them. Among these, Milgrom and Watt (2022) find market outcomes that approximate natural WE when none exist. They do not make the distinction of desiring PMEs as introduced here, nor do they explore the concept of cuts as artificial items directly. Like O’Neill et al. (2005), they have a single markup offset that can be interpreted in the language of the current paper as a single cut to the WDP-LP based on prices, similar to a tax on auction spending. The combinatorial cuts of the current paper attempt instead to generate artificial structures that are as simple or explainable as possible, in that we look for cuts reflecting simple relationships among bids that can be represented as integer knapsack constraints, with smaller constant terms favored as simpler and more explainable.
@@
Even more closely related is the recent work of Lahaie and Lubin (2025) who introduce the adaptive (endogenous) generation of price terms in a dynamic (tâtonnement) CA. They focus on CAs with unique items (supply of one item for each commodity) and “monomial” price terms that each correspond to a subset of items, allowing for the pricing of certain synergies directly, with the goal of terminating at WE prices. From the perspective of the current paper, each monomial price term can be viewed as a clique cut (i.e., a fictional item for which there is exactly one copy), though it can also be shown that not all clique cuts are implicitly considered. (For example, Example 4.1 below uses a clique cut that does not correspond to a unique subset or monomial term in that paper.) They are able to show computationally that these cuts are highly effective in generating WEs by augmenting the space of prices under consideration, complementing the perspectives offered here. Here, we also show a framework that adds pricing terms, but move to more general multi-unit CAs and explicitly search over a broader set of possible cuts. This richer set of cuts allows us to guarantee the stricter PME refinement of WE.
@@
All AWE (and hence PME) prices naturally produce core-selecting auctions, in which no coalition of bidders could propose a mutually (weakly) preferred outcome with higher total revenue. Core-selecting auctions are also described in Day and Milgrom (2008), with the precise selection of core points discussed in Day and Cramton (2012) and Bünz et al. (2022). We use the ideas of quadratic minimization as a tie-breaking rule as in the latter papers, but emphasize that here, quadratic tie-breaking is over item prices instead of personalized payments, an arguably better approach. Here we also show that all core-selecting points can be reached as an AWE via just one valid cut, which allows us to easily select the core point we want: PME prices with a minimal amount of artificial price adjustment, then decomposed into a collection of more interpretable cuts.
@@ 3 Model Preliminaries
In a combinatorial auction (CA), a single seller offers $J$ different types of indivisible items for sale to a set of $I$ bidders, denoted $\mathcal{I}=\{1,2,...i,...I\}$. The set of item types is denoted $\mathcal{J}=\{1,2,...j,...J\}$, and each item $j$ has a supply of $c_{j}\geq 1$ identical copies available to be purchased. CAs with unique items, i.e., those with $c_{j}=1$ for all $j$, are often studied in the literature; here we allow for the more general case. Column vector $c$ will denote the complete supply vector with elements $c_{j}$.
@@
All proofs not directly stated in the text are deferred to Appendix 9.
@@ 3.1 Winner Determination
The linear relaxation of WDP replaces (4) with $\bm{0}\leq x\leq\bm{1}$, and will be called WDP-LP. Because a solution to WDP-LP is potentially fractional, we write $w^{f}(\mathcal{I},\mathcal{J},\mathcal{K})$ for its objective function. Similarly, $x^{f}$ is a solution to WDP-LP, used when it is different from $x^{*}$. A valid inequality or cut in the current context is a new row added to WDP that is satisfied by all binary feasible solutions to WDP, but not necessarily all fractional feasible solutions to the relaxation WDP-LP. Because the final cuts we prescribe will have non-negative integer coefficients, the addition of any such cut results in a new WDP formulation equivalent to an extended CA with additional items but the same set of solutions, hence our interpretation of cuts as artificial items.
@@ 3.2 Payments, Prices, and Equilibrium
Definition 3.1
An outcome is an artificial Walrasian equilibrium (AWE) for a CA if and only if it is a WE for an augmented CA in which valid cuts of the original CA are treated like natural items with standard linear prices.
@@ 3.3 Computing Dual Prices
In this paper, we explore finding AWEs and will motivate and demonstrate how to select from among them over the next several pages. Anticipating the final outcome of this process (before we have fully specified it), assume that $\bm{A}$ contains a final set of added cuts that result in an AWE. It follows that WDP-LP will solve to integer optimality, and that dual prices form a WE via the main result of Bikhchandani and Mamer (1997). The dual of WDP-LP is:
@@ 4.1 Motivation and examples
which combine to yield $p_{A}+p_{B}+p_{C}\geq 15$, a violation of IR with a payment of $15>12$, the value to the winning bidder. It is easy to show that the associated WDP has a valid cut:
@@ 4.2 Price-Match Equilibria for CAs
Adomavicius and Gupta (2005) define a winning level of a bundle $a$ at any instant in a continuous efficient CA as the amount needed for a new bid on $a$ to just become winning, given the bids submitted up until that point in time. They show that it can be computed in our notation (before the new bid has been added to $\mathcal{K}$) as $w(\mathcal{I},\mathcal{J},\mathcal{K})-w(\mathcal{I},\mathcal{J}\setminus a,\mathcal{K})$, where $\mathcal{J}\setminus a$ represents the same item classes but with supply vector reduced to cac-a.
@@
Indeed, in Example 4.1, without the cut, the first three bids would be “half accepted” by solution $x^{f}=[0.5,0.5,0.5,0]^{T}$ to WDP-LP. In doing so, it’s as if the three bids are allowed to represent $15$ units of value, when really they only generate $10$ units of value integrally. In Example 4.2 (if a cut is not applied), the raised bid “hides” this non-convex over-influence of these three bids, allowing them to drive prices up beyond their ability to match those prices in any integral solution.
To further refine the winning-level notion for the current context (i.e., a multi-unit, sealed-bid CA with linear prices and multi-minded bidders), consider the following definition. As is typical, we use a i-i superscript to denote a situation in which bidder $i$ is removed or zeroed out while all other bidders remain.
@@
Directly from the definition of PME, the second-price auction for a single item has a PME outcome. Similarly, the Generalized Second-Price Auction, in which the winner of the $n$th slot pays the amount of the n+1n+1st bid (famous for its use in Google’s ad-word auctions), prescribes a PME outcome: if any bidder were to leave, each bidder in a worse-ranked slot could be moved up one slot without violating IR. Leonard (1983) shows that when unit-demand holds, VCG is the unique minimal NWE. Together with Corollary 6.6 below, we have:
@@
Unit demand is sufficient to guarantee min-NWE = PME = MRC = VCG.
@@
PMEs are a (sometimes strict) subset of the AWEs associated with a CA. Examples below show that for the same CA instance, several PMEs may be available, with some relying more heavily on artificial item prices than others, leading us to the following two concepts.
Definition 4.9
The total artificiality of an AWE is $\sum_{\mathcal{J}^{+}}\sum_{\mathcal{K}}p_{j}a_{jk}$, where $\mathcal{J}^{+}$ is the set of all added artificial items.
Definition 4.10
A minimally artificial PME (MAP) is a PME for which total artificiality is minimal among all PMEs.
Note that total artificiality includes all bids of all bidders (not just winning bids) in its summation. MAP outcomes will be shown to always exist for a CA and may not be unique. Directly from the definition, when a NWE exists, all NWE prices are MAP outcomes with zero total artificiality.
@@ 5 Comparison to Min-Revenue Core-Selecting and VCG Mechanisms
The proof sketch is that core-selecting constraints as in Equation 9 can be derived by simply adding the relevant envy-freeness constraints (5) indexed by winning bids in the $w(\mathcal{C},\mathcal{J},\mathcal{K^{\mathcal{C}}})$ outcome. (This result is well known in various forms. See, e.g., Bikhchandani and Ostroy 2002.) The converse does not directly hold: not all core outcomes can be reached via the currently available items prices. (This will be addressed later by Theorem 6.2, however.)
To demonstrate this important difference between MAP prices and the VCG and MRC benchmarks from the literature, consider the following numerical example, adapted from Beck (2013):
@@
[…]
@@
$\rho^{MRC}_{1}=5\text{ for A}$ $\rho^{MRC}_{2}=5\text{ for B}$ $\text{when selling }\{A,B\}$
$\rho^{MRC}_{1}=10\text{ for C}$ $\text{when selling }\{C\}$
$\rho^{MRC}_{1}=10\text{ for AC}$ $\rho^{MRC}_{2}=0\text{ for B}$ $\text{when selling }\{A,B,C\}$
and observe that total revenue goes down drastically (from $20$ to $10$) when combining the $\{A,B\}$ auction with that of an “unrelated good” $\{C\}$ into a single auction. Beck (2013) notes that this shows a similar “low-revenue” problem as VCG, observing that total payments collected by VCG are $10$ whether these two auctions are conducted separately or together. With this example as motivation, we observe the unrelated goods problem vanishes for MAP prices as proposed here.
@@ 6 Generating Cuts for WE and PME
We first consider generating a single valid cut. Let its coefficients be $\alpha_{k}$ for each $k\in\mathcal{K}$, collectively the vector $\alpha$, and its right-hand side be $\alpha_{0}$:
@@
which must hold for any WDP-feasible allocation $x\in\{0,1\}^{K}$. These $\alpha$ terms will be variables during the cut generation process and then later treated as constants in the final price determination, functioning exactly as all other items in $\mathcal{J}$, and concatenated as a new row of the WDP matrix $\bm{A}$. This new cut will thus be treated as an artificial item in the auction. Eventually, we prefer a collection of artificial items with each $(\alpha,\alpha_{0})\in\mathbb{Z}^{K+1}_{\geq 0}$ for interpretability, but for intermediate steps in our reasoning and algorithms we will also consider “unscaled” or continuous valid cuts, which we denote with $(\phi,\phi_{0})\in\mathbb{R}^{K+1}_{\geq 0}$.
Any newly added artificial item must be fully demanded at our fixed WDP solution $x^{*}$, i.e., $\alpha x^{*}=\alpha_{0}$ or $\phi x^{*}=\phi_{0}$. That is every copy of an artificial item must be purchased by the winning bids. If not, WE fails by definition, and pricing an excess-supply item may violate the core property, even when envy-freeness constraints (5) are satisfied. Because the preponderance of previous work on cut generation was used to solve IPs like WDP, this assumption that we know $x^{*}$ and constrain that $\alpha x^{*}=\alpha_{0}$ when looking for cuts is peculiar to the current line of research, to our knowledge.